OpenStudy (anonymous):

A rectangular pyramid has a base length, l, of 13.8ft, a base width, w, of 16.7ft, and a slant height, l, of 13.9ft. Find the surface area of the pyramid to the nearest square unit. A. 654 ft2 B. 700 ft2 C. 720 ft2 D. 800 ft2 A rectangular pyramid has a base length, l, of 13.8ft, a base width, w, of 16.7ft, and a slant height, l, of 13.9ft. Find the surface area of the pyramid to the nearest square unit. A. 654 ft2 B. 700 ft2 C. 720 ft2 D. 800 ft2 @Mathematics

OpenStudy (anonymous):

To find the area of the base use Area=length*width, so 13.8x16.7=230.46 To find the area of the sides use Area=2*(1/2)*(1/2)LH and Area=2*(1/2)*(1/2)WH and then multiple that by 2, so 2*(1/2)*(1/2)*13.8*13.9= 95.91*2=191.82 2*(1/2)*(1/2)*16.7*13.9=116.065*2=232.13 Add them all together, 191.82+232.13+230.46=654.41 Your answer is A

OpenStudy (anonymous):

area of base = 13.8 * 16.7 = 230.46 sq ft area of slanted triangles smallest = (sqrt(13.9^2 - (13.8/2)^2) * 13.8/2 ) * 2 = 191.82 largest = (sqrt(13.9^2 - (16.7/2)^2) * 16.7/2 ) * 2 = 232.13